SRC Modula-3 Commercial License

SRC Modula-3, as described in Attachment A, is distributed by Digital Equipment Corporation (“DIGITAL”), a corporation of the Commonwealth of Massachusetts. DIGITAL hereby grants to you a non-transferable, non-exclusive, royalty free worldwide license to use, copy, modify, prepare integrated and derivative works of and distribute SRC Modula-3 for commercial purposes, subject to your agreement to the following terms and conditions:

• The SRC Modula-3 Non-commercial License, Attachment B, shall be included in the code and must be retained in all copies of SRC Modula-3 (full or partial; original, modified, derivative, or otherwise):

• You acquire no ownership right, title, or interest in SRC Modula-3 except as provided herein.

• You agree to make available to DIGITAL all improvements, enhancements, extensions, and modifications to SRC Modula-3 which are made by you or your sublicensees and distributed to others and hereby grant to DIGITAL an irrevocable, fully paid, worldwide, and non-exclusive license under your intellectual property rights, including patent and copyright, to use and sublicense, without limititation, these modifications.

• SRC Modula-3 is a research work which provided “as is”, and DIGITAL disclaims all warranties with regard to this software, including all implied warranties of merchantability and fitness of purpose. In no event shall DIGITAL be liable for any special, direct, indirect, or consequential damages or any damages whatsoever resulting from loss of use, data or profits, whether in an action of contract, negligence or other tortious action, arising out of or in connection with the use or performance of this software.

• You may distribute in any manner derivative or integrated software that contains portions of the SRC Modula-3 core and runtime libraries, but not the compiler; distribute SRC Modula-3 to third parties, provided such third parties sign a letter incorporating terms substantially similar to this License and such third parties agree to abide by the terms and conditions under which the software was distributed. Further distribution by such third parties is allowed provided that the conditions set forth in this License are met and the non-commercial licence is included.

• You agree to defend, indemnify and save DIGITAL harmless against all liability, attorney’s fees and costs of defending against third party claims against DIGITAL arising out of your use, sale, license, or sublicense of SRC Modula-3.

• You may not use the name of DIGITAL in any advertisement, press release or for other publicity with reference to SRC Modula-3 without prior written consent of DIGITAL.

• You sign and return this agreement to DIGITAL at:

Systems Research Center SRC Modula-3 Licensing 130 Lytton Avenue Palo Alto, CA 94301

AGREED:

Licensee

Signature  Address

Printed Name  City, State, Zip

, 19

Title  Date

ATTACHMENT A

SRC Modula-3 includes: Compiler:

a driver program to invoke the compiler passes and a modula-3 to c compiler.

Runtime library:

a runtime kernel, a garbage collector, a threads interface and implementation, an operating system interface and implementation, a text string interface and implementation, a time interface and implementation, an unsigned word interface and implementation.

Core library:

implementations and interfaces to access files, fingerprint text strings, format basic types, scan literals of the basic types, build geometric primitives, build and maintain lists, parse command line parameters, generate random numbers, read from input streams, write to output streams, keep simple statistics, build and maintain tables, and generate unique identifiers.

Miscellaneous tools:

a pretty printer and linker.

Documentation:

the “modula-3 Report (revised),” a user’s manual, a licensing agreement, and Makefiles to construct the software.

ATTACHMENT B

SRC Modula-3 Non-commercial License

SRC Modula-3 is distributed by Digital Equipment Corporation (“DIGITAL”), a corporation of the Commonwealth of Massachusetts. DIGITAL hereby grants to you a non-transferable, nonexclusive, royalty free worldwide license to use, copy, modify, prepare integrated and derivative works of and distribute SRC Modula-3 for non-commercial purposes, subject to your agreement to the following terms and conditions:

• The SRC Modula-3 Non-commercial License shall be included in the code and must be retained in all copies of SRC Modula-3 (full or partial; original, modified, derivative, or otherwise):

• You acquire no ownership right, title, or interest in SRC Modula-3 except as provided herein.

• You agree to make available to DIGITAL all improvements, enhancements, extensions, and modifications to SRC Modula-3 which are made by you or your sublicensees and distributed to others and hereby grant to DIGITAL an irrevocable, fully paid, worldwide, and non-exclusive license under your intellectual property rights, including patent and copyright, to use and sublicense, without limititation, these modifications.

• SRC Modula-3 is a research work which is provided “as is”, and DIGITAL disclaims all warranties with regard to this software, including all implied warranties of merchantability and fitness of purpose. In no event shall DIGITAL be liable for any special, direct, indirect, or consequential damages or any damages whatsoever resulting from loss of use, data or profits, whether in an action of contract, negligence or other tortious action, arising out of or in connection with the use or performance of this software.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1990 Digital Equipment Corporation

All Rights Reserved